UT Feb 15 2020 - Split Night LBTI (1st half) - DEC (2nd half)



Staff:

  • OSA: J.Williams
  • ISp: Mtn Staff
  • Mtn Manager: B. Waggoner
  • ISA/Observer: J.Power
  • AO: Jared

Night Info:

  • Sunset 18:04
  • Nautical Twilight Ends 18:59
  • Astro Twilight Ends 19:27
  • Night Center 00:33
  • Astro Twilight Begins 05:40
  • Nautical Twilight Begins 06:08
  • Moon sets 06:51
  • Sunrise 07:03

Night Plan:

  • Starting Config: LBTI
  • DONELBTI science twilight-00:33
  • DONELBTI Engineering 00:33-03:33
  • DONEReconfigure from LBTI to MODS 03:34-03:49
  • DONEMODS Engineering 03:50-05:10
    • DONEMODS - run zernike_collect to measure rotator offset angle for active optics -1.5hr
    • DONEIf clear - collect spectrophotometric standard in all settings (Feige 34 is up)

Backup

  • If MODS and/or zernike_collect cannot be executed:
    • LUCI-Bino SL - GRB 211211A - 1.25hr (2nd half) - Rothberg
      • OT File: /home/lbto/Service_Observing/2022A_D_E/LUCI/GRB211211A/GRB211211A_LD2021B-002_V5_OT.xml
      • Ks and J (fraternal) imaging
      • Seeing < 1.3"
      • Thin Clouds
      • If clear (photometric) - execute standard GSPC P272-D before and after observations (just rotates through Ks and J)
  • Poor Weather (seeing > 2" and/or thick clouds or if MODS/zernike_collect fails and conditions too poor for GRB)
    • Local Mergers - Rothberg
    • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V40.xml
      • UGC 9829 - 2nd half
      • UGC 10607 -2nd half
      • UGC 10675 -2nd half
      • NGC 4004 -2nd half
      • Arp 156 -2nd half
      • NGC 3597 -00:00-03:30hr
    • Spectral Flat monitoring for AZ - clear only and in twilight - any seeing
    • LUCI Photometric Standards (only if clear and terrible seeing, if guider > 3" then change ACTIVE to TRACK mode in scripts - fuzzy/blurry are OK)
      • LUCI - /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards/LUCI_Phot_ZP_OT_V3.xml

Overview

0:00 Instruments started up, MODS broup up and simSnap run. Dave loaded updated LUCI flexcomp tables with data from December

LBTI Engineering

7:33 Slewing to HIP56147 for LBTI engineering. Running through all Lmircam filters with nod pairs, adjusting integration times and detector modes to acheive 50% saturation ( flux calbration). Testing PK50 blocker configurations. See LBTI night log for results

7:35 OCAM reported negative counts and required rebinning fixed. OCAM has been flakey tonight.

Completed collecting data for filters: StdM, StdL, Kshort, H (with and without PK50), MK-J (with and without PK50), N03946-4 in fast and medium, NB3950-035, NB355-041, NB3375-025, NB2925-055. Found that Kshort and MK-J require the PK50 but H does not.

Many issues experienced by LBTI:
  • Safeskipping while offsetting. The only recovery is to resent the preset. Conditions were good. Clear with good seeing. Safeskips were not noted until we offset. GOpt or NCPA related? Timestamps: 9:28, 9:49, 10:13
  • OCAM not live, coupled with Shell RIP. Cause unknown. 6:09, 8:51, 10:21
Experienced many issues first half as well. Winds were higher and thought perhaps a contributor.
  • OCAM not live: 4:55,
  • OCAM not live coupled with shell rip 3:58, 6:09
  • OCAM shows negative counts, correction lost. Rebin fixes. 7:04
  • Safekipping during offset 4:39

zernike_collect for MODS/MODS

10:43 Reconfig to MODS for zernike coefficient work

11:00 Test preset to 12.75 mag star. This is the star to be used for the zernike work. Josh pointed and collimated.

11:10 Set up idl. Stopped WFS

11:13 Started zernike_collect on both sides, using the z5_z10 list. Working well, zernikes are changing as expected. Skies are clear with seeing 1"
left_wfs000001-25
right_wfs000001-25

11:23 Starting zernike_collect on both sides, using the z14_z15 list.

left_wfs000026-34
right_wfs000026-34

Did I miss something? No. It is quick when everything works.

wfsc data .fits files, and analysis logs are in /home/telescope/data/20220215/zernike_ldg and zernike_rdg respectively.

JMH (the next afternoon): I've adjusted WFS_rotAngleOffs from 120 deg to 117 deg for both MODS AGWs. This is the offset of the angle that the measured Zernikes are rotated to get into primary mirror coordinates. It is a small change, but should make MODS active optics 1% more stable.

P.S. Last night's data seemed to also confirm that there is a problem with the reconstructor (or the reading of the reconstructor) for Z14/Z15. The problem is similar for MODS and PEPSIPOL with different AGWs. Watch for a new IT coming to a theater near you.

Spectrophotometric standards for MODS/MODS

11:29 Slewing to specphot HZ44

MODS1:

Computed Slit Alignment Offset:
 dX = 0.499 arcsec
 dY = 9.970 arcsec
MODS1 Offset Command:
 offsetxy 0.499 9.970 rel

MODS2:

Computed Slit Alignment Offset:
 dX = 5.433 arcsec
 dY = 7.808 arcsec
MODS2 Offset Command:
 offsetxy 5.433 7.808 rel

MODS2 additional offset:

MODS2 Offset Command:
offsetxy -0.713 0.144 rel

11:38 Starting science on HZ44

mods1b.20220215.0003-14, mods1r.20220215.0005-16
mods2b.20220215.0003-14, mods2r.20220215.0006-17

12:36 Sending preset to BD33

MODS2:

Computed Slit Alignment Offset:
 dX = 4.353 arcsec
 dY = 7.780 arcsec
MODS2 Offset Command:
 offsetxy 4.353 7.780 rel

Additional offset for mods2:

Computed Slit Alignment Offset:
 dX = -0.429 arcsec
 dY = 0.331 arcsec
MODS2 Offset Command:
 offsetxy -0.429 0.331 rel

MODS1 readout lag, The offsets did not appear int eh terminal but sent -0.358, 10.772

12:45 Starting science on bd33. Twilight is approaching. This target is transiting in ~50min but the sun will get us before then. Will execute as many configs as is reasonable before the sun becomes and issue. Seeing has deteriorated, up to 1.8" now.

13:11 Completed the Dual grating and red grating for BD33. 12 deg Twilight now so closing up

Calibrations:

mods: 0.5" dual grating flats, all bias


-- BarryRothberg - 11 Feb 2022
Topic revision: r8 - 15 Feb 2022, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback